OpenStudy (anonymous):

How do you solve 2cosØ=-1 for 0<Ø<360degrees?

OpenStudy (anonymous):

Can you help with another one? @CGGURUMANJUNATH

OpenStudy (anonymous):

Can anyone help please?!

OpenStudy (anonymous):

Can you help @CGGURUMANJUNATH?

OpenStudy (anonymous):

start with \[\cos(\theta)=-\frac{1}{2}\]

OpenStudy (anonymous):

Divide each side by 2. You get cos x = -1/2. Look on the unit circle; which value has a cosine of -1/2?

OpenStudy (anonymous):

2pi/3

OpenStudy (anonymous):

then see if you can think of an angle or two that has that value for cosine as @aurorablue01 said, look at the unit circle the last page of the attached cheat sheet will help

OpenStudy (anonymous):

and 4pi/3?

OpenStudy (anonymous):

apparently you are supposed to give your answer in degrees there are two answers in the interval

OpenStudy (anonymous):

Oh, so it is 120 and 240?

OpenStudy (anonymous):

yeah your two are right

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok, thanks both of you.

OpenStudy (anonymous):

You're welcome